OpenStudy (anonymous):

Refer to the graph of y = sin x or y = cos x to find the exact values of x in the interval [0, 4π] that satisfy the equation for 2cos x = −2

OpenStudy (anonymous):

\[2\cos(x)=-1\iff \cos(x)=-1\]

OpenStudy (anonymous):

cosine is \(-1\) at \(x=\pi, 3\pi,5\pi,...\)

OpenStudy (anonymous):

It says that is not right... Which is why I am confused.

OpenStudy (anonymous):

in the interval \([0,4\pi]\) the two answers are \(\pi\) and \(3\pi\)

OpenStudy (anonymous):

oh ok that makes sense. Yes that is right then.

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

I appreciate it thanks.